hello, i am having a hard time understanding how to solve this word problem. Could you show me a way of answering this problem? please and thank you
a merchant increases the wholesale cost of a washing machine by 30% to determine the selling price. If the washer sells for $588.90, find the wholesale cost.

Answer by Maths68(1474)

Let
Original wholesale cost = x
30% of x = 30/100 (x)
30% of x = 3/10 (x)
30% of x = 3x/10
original cost + 30% of original cost = Selling price
x+3x/10=588.90
multiply by 10 both sides of the equation
10(x+3x/10)=588.90*10
10x+3x=5889
13x=5889
Divide by 13 both sides of the equation
13x/13=5889/13
x=453
Original wholesale cost = 453$